# The tame Deligne-Simpson problem

### v1

@@ 1.3. Crawley-Boevey’s conjecture
Crawley-Boevey informed me that he also has an uncirculated but complete proof since May 2018.
@@ 1.4. Strategy of the proof
In a previous work of Schedler-Tirelli [ST22], the possible dimension vectors of simple representations beyond the set $\Sigma_{\mathbf{q}}$ have been limited to types close to affine Dynkin (necessarily simply-laced). This is analogous to the results of Crawley-Boevey for additive quiver varieties; see [CB01, §8], where he used some hard algebra to rule out these particular cases. However, his method does not seem to have an obvious multiplicative counterpart, and therefore we propose to turn to geometric tools to circumvent the difficulty.
@@ Acknowledgement
I would like to thank Pengfei Huang, Emmanuel Letellier and Xueqing Wen for some helpful discussions. I also thank William Crawley-Boevey for kindly informing me of his results.
@@ 2.2. Schedler-Tirelli’s classification of dimension vectors of simple representations
    Isotropic imaginary roots. Note that the support of an isotropic root is necessarily an affine Dynkin diagram. Let $Q$ be a star-shaped affine Dynkin quiver and $\boldsymbol{\delta}$ the minimal positive imaginary root. Then, $\mathbf{d}=m\boldsymbol{\delta}$ for some $m\in\mathbb{Z}_{>0}$.
@@
([ST22, Corollary 6.18]) Suppose that there exists a simple representation in $\mathcal{M}(\mathbf{q},\mathbf{d})$. Then, one of the following occurs:
@@
    $\mathbf{d}=e_{\infty}+m\boldsymbol{\delta}$ is a flat root and $m\geq 2$. Moreover, we have $q_{\infty}=\mathbf{q}^{\boldsymbol{\delta}}=1$.
@@ 2.5. Reduction to semi-simple character varieties
(2.5.1) $\mathcal{M}(\mathbf{q},\mathbf{d})$$\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},\mathbf{d})$$\mathcal{M}_{B}(\overline{\mathcal{C}})$$\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi})$$\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})$$\mathcal{M}_{B}(\mathcal{C}^{\prime})$$\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi}^{\prime})$$\mathcal{M}_{Dol}(\mathbf{d},\boldsymbol{\alpha},(\mathcal{O}_{j})_{j}).$$\scriptstyle{\mathbf{Iso}_{1}}$$\scriptstyle{\mathbf{Var}_{0}}$$\scriptstyle{\mathbf{Iso}_{2}}$$\scriptstyle{\mathbf{Var}_{1}}$$\scriptstyle{\mathbf{NH}_{1}}$$\scriptstyle{\mathbf{Iso}_{3}}$$\scriptstyle{\mathbf{NH}_{2}}$$\scriptstyle{\mathbf{Var}_{2}}$
@@
  • $\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi})$ is a moduli space of filtered local systems. In the associated graded of the filtered structures, the monodromies are scalar matrices; however, the same scalar matrix may appear in different graded spaces.
  • $\mathcal{M}_{B}(\mathbf{d}^{\prime},\boldsymbol{\beta}^{\prime},\boldsymbol{\xi}^{\prime})$ is another moduli space of filtered local systems. The monodromies are again scalar matrices in the associated graded, but the monodromies can distinguish different graded spaces.
@@
  • $\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})$ is a moduli space of parabolic Higgs bundles. The residues of the Higgs fields lie in prescribed semi-simple adjoint orbits when passing to the Levi quotients of the parabolic structures, but these adjoint orbits need not be central in each factor of the Levi.
  • $\mathcal{M}_{Dol}(\mathbf{d},\boldsymbol{\alpha},(\mathcal{O}_{j})_{j})$ is another moduli space of parabolic Higgs bundles. The residues of the Higgs fields are central when passing to the Levi quotients.
@@
  • $\mathbf{Iso}_{1}$ and $\mathbf{Iso}_{2}$ are the identification between multiplicative quiver varieties and the moduli spaces of (filtered) local systems. More precisely, $\mathbf{Iso}_{1}$ is the isomorphism (1.6), and $\mathbf{Iso}_{2}$ is given by [Yam08, Theorem 1.2] where we need $\theta_{v}>0$ for all $v\neq\star$.
@@ 2.6. The proof of Theorem A
The case $(\mathbf{Aff})$. Regard $\rho^{\prime}_{s}$ as an element of $\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi})$. Then, $\mathbf{NH}_{1}$ sends it to a $\boldsymbol{\alpha}^{\prime}$-stable parabolic Higgs bundle $(E^{\prime}_{s},\Phi^{\prime}_{s})$. We will show in Proposition 3.7 that there exists an $\boldsymbol{\alpha}$-stable parabolic Higgs bundle $(E_{s},\Phi_{s})$ that is mapped to $(E^{\prime}_{s},\Phi^{\prime}_{s})$ by $\mathbf{Var}_{2}$; here, $\boldsymbol{\alpha}$ is chosen to be almost generic (see Definition 3.3). Passing to the Betti side of $\mathbf{NH}_{2}$, we find a $\boldsymbol{\beta}$-stable filtered local system $\mathscr{L}$, where the eigenvalues $\boldsymbol{\xi}^{\prime}$ are almost generic. Then, Proposition 3.11 will show that $\mathbf{Iso}_{3}(\mathscr{L})$ is an irreducible local system with monodromies in $\mathcal{C}^{\prime}$. However, this contradicts Kostov’s Theorem 2.7. We conclude that there exists no simple representation in $\mathcal{M}(\mathbf{q},\mathbf{d})$.
@@
Lemma 2.8.
Suppose that the quiver, $\mathbf{d}$ and $\mathbf{q}$ are of type $(\mathbf{Aff}^{\infty})$. Then, we have $\dim\mathcal{M}(\mathbf{q},\mathbf{d})=2p(\mathbf{d})=2m$.
Proof.
@@
This is in fact the defining property of flat roots in [ST22], and it is a part of the statement of [ST22, Theorem 6.16]. (The proof of this theorem is based [Su06], and what we actually need here is [Su06, Proposition 4.2].) Combined with [CBS06, Lemma 6.2, Corollary 7.3], this implies that $\mu^{-1}(\mathbf{q})$ is equidimensional of dimension $2p(\mathbf{d})+\dim G-1$, where $G=\prod_{v\in Q_{0}}\operatorname{GL}_{d_{v}}$. It follows that $\dim\mathcal{M}(\mathbf{q},\mathbf{d})$ has dimension at least $2p(\mathbf{d})$. However, the inequality (2.6.1) and [CBS06, Lemma 7.1] imply that the dimension is at most $2p(\mathbf{d})$. ∎
By [CBS06, Lemma 7.1], the locus of semi-simple representations of type $(k_{i},\mathbf{d}^{(i)})_{1\leq i\leq r}$, if nonempty, has dimension $\sum_{i=1}^{r}2p(\mathbf{d}^{(i)})$; in particular, the locus of simple representations has dimension $2p(\mathbf{d})=2m$, and the locus consisting of direct sums of distinct $\boldsymbol{\delta}$-dimensional simple representations also has dimension 2m. However, the latter stratum is nonempty in view of Crawley-Boevey-Shaw’s theorem [CBS06, Theorem 1.1]. We conclude that $\mathbf{d}$-dimensional simple representations do not exist.
@@ 3.1. Combinatorial data
$\boldsymbol{\tau}(\underline{\nu}):=\{[j,i]|1\leq j\leq k,\leavevmode\nobreak\ 0\leq i\leq\nu_{j}\}$
where $[j,i]$ means a pair consisting of integers $j$ and $i$ (instead of an interval); we will call $\boldsymbol{\tau}(\underline{\nu})$ a type, which will be associated to a filtered structure or a parabolic structure. An element of $\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$ will be written as $\mathbf{d}=(d_{[j,i]})_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}$ with $d_{[j,i]}\in\mathbb{Z}_{\geq 0}$, and similarly $\mathbf{q}=(q_{[j,i]})_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}\in(\mathbb{C}^{\ast})^{\boldsymbol{\tau}(\underline{\nu})}$. We define a star-shaped quiver $Q=(Q_{0},Q_{1})$ from $\boldsymbol{\tau}(\underline{\nu})$ in the following way. Let $Q^{\prime}_{0}=\boldsymbol{\tau}(\underline{\nu})$ and let $Q^{\prime}_{1}$ be the set of arrows $a_{[j,i]}:[j,i]\rightarrow[j,i+1]$ for $j,i]\in\boldsymbol{\tau}(\underline{\nu})$ and $i\neq\nu_{j}$. We define $Q_{0}$ as $Q^{\prime}_{0}$ modulo the relation $j,0]\sim[j^{\prime},0]$ for any $1\leq j\leq j^{\prime}\leq k$ and define $Q_{1}$ as the induced set of arrows among the elements of $Q_{0}$. We will say that $\mathbf{d}$ has rank $n$ if $d_{[j,0]}=n$ for all $j$; we may regard such a $\mathbf{d}$ as an element of $\mathbb{Z}_{\geq 0}^{Q_{0}}$. Conversely, any $\mathbf{d}\in\mathbb{Z}_{\geq 0}^{Q_{0}}$ can be regarded as a rank $n$ vector in $\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$.
@@
that preserves the $j$-component while being increasing in the $i$-component and satisfies $\sigma([j,0])=[j,0]$ for every $j$. By abuse of notation, we may write $\sigma([j,i])=[j,\sigma(i)]$. For any $\mathbf{d}\in\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$, the value of $\sigma^{\ast}\mathbf{d}$ at $[j,i]$ is equal to $d_{[j,\sigma(i)]}$; this should be thought of as the dimension vector of a degenerated flag to be defined later.
Given a $\mathbf{d}\in\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$, it will be convenient to introduce the associated vector
with $d_{[j,i]}^{\ast}=d_{[j,i]}-d_{[d,i+1]}$ for all $[j,i]$.
@@ 3.4. Nonabelian Hodge theory
The nonabelian Hodge correspondence between the moduli space of filtered local systems and the moduli space of parabolic Higgs bundles was established by Simpson in [Sim90]. The original statement is a bijection between the stable objects on each side, and a homeomorphism between the entire moduli spaces is recently proved by Huang-Sun:
@@
([HS25, Theorem 5.3]) The nonabelian Hodge correspondence induces a homeomorphism of topological spaces
@@
Example 3.2.
Let us give an example to show how the more familiar nonabelian Hodge correspondence between strongly parabolic Higgs bundles and local systems with semi-simple monodromies arises as a special case. On the Dolbeault side, the eigenvalues of the residues of Higgs fields are zero, and thus we have $b=c=0$. The vanishing of $b$ means that on the Betti side the filtered structure is trivial with weight 0, whereas the vanishing of $c$ means that the eigenvalues on the Betti side are of the form $\exp(-2\pi i\alpha)$, where $\alpha$ is a parabolic weight.

### v2

@@ 1.3. Crawley-Boevey’s conjecture
There is another proof of this theorem by Crawley-Boevey in his latest preprint [CB25], which according to him has been uncirculated but complete since May 2018.
@@ 1.4. Strategy of the proof
In a previous work of Schedler-Tirelli [ST22], the possible dimension vectors of simple representations beyond the set $\Sigma_{\mathbf{q}}$ have been limited to types close to affine Dynkin (necessarily simply-laced). This is analogous to the results of Crawley-Boevey for additive quiver varieties; see [CB01, §8], where he used some hard algebra to rule out these particular cases. In a recent preprint [CB25], he has managed to rule out these cases in the multiplicative setting, and the proof seems to be more difficult than in the additive setting. In this article, we propose to turn to geometric tools to circumvent the difficulty.
@@ Acknowledgement
I would like to thank Pengfei Huang, Emmanuel Letellier and Xueqing Wen for some helpful discussions. My thanks also go to Philip Boalch, who have oriented me through the literature on nonabelian Hodge theory. I am grateful to William Crawley-Boevey for kindly informing me of his results, encouraging me to release my preprint, and especially for developing many of the tools that are used in this paper.
@@ 2.2. Schedler-Tirelli’s classification of dimension vectors of simple representations
    Isotropic imaginary roots. Note that the support of an isotropic root is necessarily an affine Dynkin diagram, say $Q$. Let $\boldsymbol{\delta}$ be the minimal positive imaginary root supported on $Q$. Then, $\mathbf{d}=m\boldsymbol{\delta}$ for some $m\in\mathbb{Z}_{>0}$.
@@
([ST22, Corollary 6.18]) Suppose that $\mathbf{d}$ is defined on a star-shaped quiver and that there exists a simple representation in $\mathcal{M}(\mathbf{q},\mathbf{d})$. Then, one of the following occurs:
@@
    $\mathbf{d}=e_{\infty}+m\boldsymbol{\delta}$ is a flat root with $m\geq 2$, where $e_{\infty}$ is the simple root corresponding to the vertex $\infty$. Moreover, we have $q_{\infty}=\mathbf{q}^{\boldsymbol{\delta}}=1$.
@@ 2.5. Reduction to semi-simple character varieties
(2.5.1) $\mathcal{M}(\mathbf{q},\mathbf{d})$$\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},\mathbf{d})$$\mathcal{M}_{B}(\overline{\mathcal{C}})$$\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi})$$\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})$$\mathcal{M}_{B}(\mathcal{C}^{\prime})$$\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi}^{\prime})$$\mathcal{M}_{Dol}(\tilde{\mathbf{d}},\boldsymbol{\alpha},(\mathcal{O}_{j})_{j}).$$\scriptstyle{\mathbf{Iso}_{1}}$$\scriptstyle{\mathbf{Var}_{0}}$$\scriptstyle{\mathbf{Iso}_{2}}$$\scriptstyle{\mathbf{Var}_{1}}$$\scriptstyle{\mathbf{NH}_{1}}$$\scriptstyle{\mathbf{Iso}_{3}}$$\scriptstyle{\mathbf{NH}_{2}}$$\scriptstyle{\mathbf{Var}_{2}}$
@@
  • $\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi})$ is a moduli space of filtered local systems. In the associated graded of the filtered structures, the monodromies are scalar matrices; however, the same eigenvalue may appear in different graded spaces.
  • $\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi}^{\prime})$ is another moduli space of filtered local systems. The monodromies are again scalar matrices in the associated graded, but the eigenvalues can distinguish different graded spaces.
@@
  • $\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})$ is a moduli space of parabolic Higgs bundles. The residues of the Higgs fields lie in prescribed semi-simple adjoint orbits when passing to the Levi quotients of the parabolic structures, but these adjoint orbits need not be central in the Levi.
  • $\mathcal{M}_{Dol}(\tilde{\mathbf{d}},\boldsymbol{\alpha},(\mathcal{O}_{j})_{j})$ is another moduli space of parabolic Higgs bundles. The parabolic structures refine those in $\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})$ in such a way that the residues of the Higgs fields are central when passing to the Levi quotients.
@@
  • $\mathbf{Iso}_{1}$ and $\mathbf{Iso}_{2}$ are the identification between multiplicative quiver varieties and the moduli spaces of (filtered) local systems. More precisely, $\mathbf{Iso}_{1}$ is the isomorphism (1.6), and $\mathbf{Iso}_{2}$ is given by [Yam08, Theorem 1.2] where we need $\theta_{v}>0$ for all $v\neq\star$. The dimension vectors on the two sides of $\mathbf{Iso}_{2}$ are the same in a suitable sense; see [Yam08, Theorem 4.16] and §3.1.
@@ 2.6. The proof of Theorem A
The case $(\mathbf{Aff})$. Regard $\rho^{\prime}_{s}$ as an element of $\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi})$. Then, $\mathbf{NH}_{1}$ sends it to a $\boldsymbol{\alpha}^{\prime}$-stable parabolic Higgs bundle $(E^{\prime}_{s},\Phi^{\prime}_{s})$. We will show in Proposition 3.8 that there exists an $\boldsymbol{\alpha}$-stable parabolic Higgs bundle $(E_{s},\Phi_{s})$ that is mapped to $(E^{\prime}_{s},\Phi^{\prime}_{s})$ by $\mathbf{Var}_{2}$; here, $\boldsymbol{\alpha}$ is chosen to be almost generic (see Definition 3.5). Passing to the Betti side of $\mathbf{NH}_{2}$, we find a $\boldsymbol{\beta}$-stable filtered local system $\mathscr{L}$, where the eigenvalues $\boldsymbol{\xi}^{\prime}$ are almost generic. Then, Proposition 3.12 will show that $\mathbf{Iso}_{3}(\mathscr{L})$ is an irreducible local system with monodromies in $\mathcal{C}^{\prime}$. However, this contradicts Kostov’s Theorem 2.7. Indeed, the conjugacy classes $\mathcal{C}^{\prime}$ satisfy the assumptions in Kostov’s theorem by Lemma 3.2, Remark 3.3 and Lemma 3.11. We conclude that there exists no simple representation in $\mathcal{M}(\mathbf{q},\mathbf{d})$.
@@
This is in fact the defining property of flat roots in [ST22], and it is a part of the statement of [ST22, Theorem 6.16]. (The proof of this theorem is based on [Su06], and what we actually need here is [Su06, Proposition 4.2].) By [CBS06, Lemma 7.1], the locus of semi-simple representations of type $(k_{i},\mathbf{d}^{(i)})_{1\leq i\leq r}$, if nonempty, has dimension $\sum_{i=1}^{r}2p(\mathbf{d}^{(i)})$. It follows from the inequality (2.6.1) that every representation type stratum has dimension at most $2p(\mathbf{d})$. The locus consisting of direct sums of distinct $\boldsymbol{\delta}$-dimensional simple representations has dimension 2m and is nonempty in view of Crawley-Boevey-Shaw’s theorem [CBS06, Theorem 1.1]. Now, the irreducibility of $\mathcal{M}(\mathbf{q},\mathbf{d})$ implies that there is only one stratum of dimension $2m$. Therefore, we have $\dim\mathcal{M}(\mathbf{q},\mathbf{d})=2p(\mathbf{d})=2m$. Since the locus of simple representations also has dimension $2p(\mathbf{d})=2m$ if it is nonempty, we conclude that $\mathbf{d}$-dimensional simple representations do not exist.
@@ 3.1. Combinatorial data
$\boldsymbol{\tau}(\underline{\nu}):=\{[j,i]|1\leq j\leq k,\penalty 10000\ 0\leq i\leq\nu_{j}\}$
where $[j,i]$ means a pair consisting of integers $j$ and $i$ (instead of an interval); we will call $\boldsymbol{\tau}(\underline{\nu})$ a type, which will be associated to a filtered structure or a parabolic structure. An element of $\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$ will be written as $\mathbf{d}=(d_{[j,i]})_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}$ with $d_{[j,i]}\in\mathbb{Z}_{\geq 0}$, and similarly $\mathbf{q}=(q_{[j,i]})_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}\in(\mathbb{C}^{\ast})^{\boldsymbol{\tau}(\underline{\nu})}$. For $\mathbf{d}\in\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$, it will be convenient to introduce the associated vector
with $d_{[j,i]}^{\ast}=d_{[j,i]}-d_{[d,i+1]}$ for all $[j,i]$. Following Crawley-Boevey, we say that $\mathbf{d}$ is strict if every component of $\mathbf{d}^{\ast}$ is nonnegative. We define a star-shaped quiver $Q=(Q_{0},Q_{1})$ from $\boldsymbol{\tau}(\underline{\nu})$ in the following way. Let $Q^{\prime}_{0}=\boldsymbol{\tau}(\underline{\nu})$ and let $Q^{\prime}_{1}$ be the set of arrows $a_{[j,i]}:[j,i]\rightarrow[j,i+1]$ for $j,i]\in\boldsymbol{\tau}(\underline{\nu})$ and $i\neq\nu_{j}$. We define $Q_{0}$ as $Q^{\prime}_{0}$ modulo the relation $j,0]\sim[j^{\prime},0]$ for any $1\leq j\leq j^{\prime}\leq k$ and define $Q_{1}$ as the induced set of arrows among the elements of $Q_{0}$. We will say that $\mathbf{d}$ has rank $n$ if $d_{[j,0]}=n$ for all $j$ and it is strict; we may regard such a $\mathbf{d}$ as an element of $\mathbb{Z}_{\geq 0}^{Q_{0}}$. Conversely, any strict $\mathbf{d}\in\mathbb{Z}_{\geq 0}^{Q_{0}}$ can be regarded as a rank $n$ vector in $\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$.
@@
that preserves the $j$-component while being increasing in the $i$-component and satisfies $\sigma([j,0])=[j,0]$ for every $j$. By abuse of notation, we may write $\sigma([j,i])=[j,\sigma(i)]$. For any $\mathbf{d}\in\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$, the value of $\sigma^{\ast}\mathbf{d}$ at $[j,i]$ is equal to $d_{[j,\sigma(i)]}$; this should be thought of as the dimension vector of a degenerated flag.
@@ 3.4. Nonabelian Hodge theory
The nonabelian Hodge correspondence between the moduli space of filtered local systems and the moduli space of parabolic Higgs bundles was established by Simpson in [Sim90], where he proved a bijection between the stable objects on each side. It is possible to extend this bijection to a homeomorphism between the full moduli spaces. For the following theorem, see Biquard-García-Prada-Mundet i Riera [BGPMiR20, Theorem 7.10] (the published version) or Huang-Sun’s recent preprint [HS25, Theorem 5.3]. Note that what is usually called the nonabelian Hodge correspondence is a map between the Dolbeault moduli space and the de Rham moduli space, and in the theorem below we have tacitly added to it the Riemann-Hilbert correspondence. As is pointed out to me by Boalch, the Riemann-Hilbert is nontrivial for all the noncompact generalisations. To simplify notations and terminologies, we will nevertheless hide the Riemann-Hilbert in the nonabelian Hodge. For the actual nonabelian Hodge correspondence, it is known since the earlier work of Biquard-Boalch [BB04, Theorem 6.1] that it induces a diffeomorphism between the stable loci of moduli spaces. Some special cases of this diffeomorphism were also proved by Konno [Kon93] and Nakajima [Nak96].
@@
The nonabelian Hodge correspondence induces a homeomorphism of topological spaces
@@
Lemma 3.2.
Consider the morphisms
$\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},\mathbf{d})\mathrel{{\mathop{\longrightarrow}\limits^{\mathbf{Iso}_{2}}}}\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi})\mathrel{{\mathop{\longrightarrow}\limits^{\mathbf{NH}_{1}}}}\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j}),$
where the objects on the Betti and Dolbeault side of $\mathbf{NH}_{1}$ have types $\boldsymbol{\tau}(\underline{\nu})$ and $\boldsymbol{\tau}(\underline{\mu})$ respectively. Suppose that $\mathbf{d}=m\boldsymbol{\gamma}$ for some indivisible $\boldsymbol{\gamma}$ and $\mathbf{q}^{\boldsymbol{\gamma}}$ is an $l$-th primitive root of unity. Let $e\in\mathbb{Z}$ be such that
$e+\sum_{[j,i]\in\boldsymbol{\tau}(\underline{\mu})}\alpha^{\prime}_{[j,i]}d^{\prime\ast}_{[j,i]}=0.$
Then, the vector $(e,\mathbf{d})$ is indivisible if and only if $l=m$.
Remark 3.3.
[…]
Proof.
Suppose that $l<m$ and write $m=lm^{\prime}$. Note that we have tacitly identified the strict $\mathbf{d}\in\mathbb{Z}_{\geq 0}^{Q_{0}}$ with a rank $n$ vector in $\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$. Obviously, $\mathbf{d}$ is divisible by $m^{\prime}$. We have
$1=\mathbf{q}^{l\boldsymbol{\gamma}}=\prod_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}\xi_{[j,i]}^{l\gamma^{\ast}_{[j,i]}}.$
If the argument of $\xi_{[j,i]}$ is $2\pi ia_{[j,i]}$, then $\sum_{[j,i]}a_{[j,i]}l\gamma^{\ast}_{[j,i]}$ is an interger. Under the nonabelian Hodge correspondence, we have
$m^{\prime}\sum_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}a_{[j,i]}l\gamma^{\ast}_{[j,i]}=\sum_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}a_{[j,i]}d^{\ast}_{[j,i]}=\sum_{[j,i]\in\boldsymbol{\tau}(\underline{\mu})}\alpha^{\prime}_{[j,i]}d^{\prime\ast}_{[j,i]},$
since each summand $\alpha^{\prime}_{[j,i]}d^{\prime\ast}_{[j,i]}$ is a sum of those $a_{[j,i^{\prime}]}d^{\ast}_{[j,i^{\prime}]}$ with $a_{[j,i^{\prime}]}=\alpha^{\prime}_{[j,i]}$. It follows that $e$ is divisible by $m^{\prime}$, and thus $(e,\mathbf{d})$ is divisible. Conversely, suppose that $(e,\mathbf{d})=m^{\prime}(e^{\prime},l\boldsymbol{\gamma})$ for some $m^{\prime}>1$. Then, $\sum_{[j,i]}a_{[j,i]}l\gamma^{\ast}_{[j,i]}$ is an integer, which implies $\mathbf{q}^{l\boldsymbol{\gamma}}=1$. ∎